If [sin-1cos-1tan-1x]=1 where [x] denotes the greatest integer function then x lies in
a) [tan sin cos1, tan sin cos sin1]
b)(tan sin cos1, tan sin cos sin1)
c)[-1,1]
d)[sin cos tan1, sin cos sin tan1]

Grade:11

1 Answers

Jitender Singh IIT Delhi
8 years ago
Ans:
[sin^{-1}cos^{-1}tan^{-1}x] = 1
1\leq sin^{-1}cos^{-1}tan^{-1}x <2
sin1\leq sinsin^{-1}cos^{-1}tan^{-1}x <sin2
sin1\leq cos^{-1}tan^{-1}x <sin2
cossin1\leq coscos^{-1}tan^{-1}x <cossin2
cossin1\leq tan^{-1}x <cossin2
tancossin1\leq tantan^{-1}x <tancossin2
tancossin1\leq x <tancossin2
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty